# Vulnerability-Aware Spatio-Temporal Learning for Generalizable and Interpretable Deepfake Video Detection

### v2

@@ 3.2 FakeSTormer
Our multi-task framework called FakeSTormer is inspired by [42, 41], where auxiliary branches are designed to push the feature extractor to focus on vulnerabilities. As discussed earlier, the vulnerability is defined in [42, 41] as the pixel/patches that are the most likely to be impacted by blending artifacts. This strategy is therefore claimed to allow the detection of subtle artifacts that are generic across different types of manipulations. While such a vulnerability-driven approach has shown very promising results [42, 41], it does not take into account the temporal nature of videos. Therefore, in addition to spatial vulnerabilities, we argue that there is a need to model temporal vulnerabilities, which we define as significant temporal changes in the blending boundary. Specifically, we introduce two additional branches, namely a temporal head $h$ and a spatial one $g$. The branch $h$ predicts the derivatives of the blending boundary over time which can reflect high changes, typically characterizing temporal artifacts. Moreover, we suggest the use of a spatial branch $g$ which enables predicting soft labels representing the forgery intensity encoded in each frame, computed from vulnerability information. The proposed framework relies on the TimeSformer backbone [2], which we revisit for better modeling spatial and temporal information.
@@ 3.2.1 Backbone: Revisited TimeSformer
For that purpose, we attach in each dimension of $\mathbf{Z}^{0}$, a spatial token $\mathbf{z}_{s}^{0}\in\mathbb{R}^{D}$ and a temporal token $\mathbf{z}_{t}^{0}\in\mathbb{R}^{D}$, respectively. These tokens will independently interact only with patch embeddings belonging to their dimension axis by leveraging the decomposed SA [2]. This mechanism not only facilitates the disentanglement learning process of spatio-temporal features but is also beneficial to optimize the computational complexity of $\mathcal{O}(T^{2}+N^{2})$ as compared to $\mathcal{O}(T^{2}\cdot N^{2})$ in vanilla SA. Those tokens will be then fed into $L$ ($L=12$ as default) transformer encoder blocks, as described in Figure 2. Formally, the feature extraction process can be summarized as follows,
@@ 3.2.2 Temporal Head $h$
Ground Truths. Our temporal head $h$ aims to model fine-grained temporal vulnerabilities in deepfake videos through a regression task. First, to generate ground truth data for the branch $h$, we hypothesize that temporal high-changes in the blending boundary can reflect the presence of temporal artifacts (see Figure 4). To achieve this, we compute $\mathbf{D}$ based on $\tilde{\mathbf{B}}$ such that:
@@
Objective Function. For training the temporal branch, we optimize the following Mean Square Error (MSE) loss,
@@ 3.2.3 Spatial Head $g$
Ground Truths. To avoid overfitting one type of artifact, we enforce the model to explicitly predict soft labels representing the intensity level of spatial artifacts for each video frame. Note that, several works [4, 51, 48, 61, 40] have leveraged soft labels for training regularization. Given a pseudo-fake video $\mathbf{X}=(\mathbf{X}(t))_{t\in[[1,T]]}$ formed by $T$ frames and $\tilde{\mathbf{B}}=(\tilde{\mathbf{B}}(t))_{t\in[[1,T]]}$ its associated cutout blending boundary, the ground truth for these soft labels is generated for each frame $t$ as follows,
@@ 3.2.4 Overall Training Objective
Method Inter. Training Test set AUC (%)
Fake CDF DFD DFDCP DFDC DFW
Xception [45] $\times$ 73.7 - - 70.9 -
MATT [64] $\times$ 68.3 92.9 63.0 - 65.7
RECCE [3] $\times$ 70.9 98.2 - - 68.2
SFDG [54] $\times$ 75.8 88.0 73.6 - 69.3
LSDA [59] $\times$ 91.1 - 77.0 - -
STIL [22] $\times$ 75.6 - - - -
LipForensics [24] $\times$ 82.4 - - 73.5 -
RealForensics [25] $\times$ 86.9 82.2 75.9 - -
FTCN [65] $\times$ 86.9 94.4 74.0 71.0 -
ISTVT [62] $\times$ 84.1 - 74.2 - -
AltFreezing [55] $\times$ 89.5 98.5 - - -
Swin+TALL [56] $\times$ 90.8 - 76.8 - -
StyleLatentFlows [9] $\times$ 89.0 96.1 - - -
FakeSTormer ($T=4$) $\times$ 92.4 98.5 90.0 74.6 74.2
FakeSTormer ($T=8$) $\times$ 92.4 98.2 90.0 74.9 75.9
FakeSTormer ($T=16$) $\times$ 92.8 98.6 90.2 75.1 75.3
Table 1: Generalization to unseen datasets. AUC (%) comparisons on multiple unseen datasets [35, 16, 13, 14, 68]. All detectors are trained on FF++ (c23). Results are directly extracted from the original papers and from [42, 24]. Bold and Underlined text respectively highlight the best and the second best performance, excluding the variants of our framework with $T=8$ and $T=16$. Inter. denotes interpretability.
@@ 4 Experiment
Method Inter. Training set Test set AUC (%)
Real Fake CDF DFDCP
Face X-ray [32] 79.5 -
PCL+I2G [63] $\times$ 90.0 74.3
SLADD [6] 79.7 -
SBI [47] $\times$ $\times$ 93.2 86.2
LAA-Net [42] $\times$ 95.4 86.9
STC-Scratch [33] $\times$ $\times$ 83.4 86.8
STC-Pretrain [33] $\times$ $\times$ 95.8 89.4
ST-SBV [23] $\times$ $\times$ 90.3 91.2
FakeSTormer $\times$ 96.5 94.1
Table 2: Comparison with other data synthesis methods. For fair comparison, we train our FakeSTormer on raw data of FF++ (c0) and test on CDF [35] and DFDCP [13].
@@ 4.1 Settings
Datasets. We set up our datasets following several works [55, 65, 56, 62, 9, 58, 54, 3]. For both training and validation, we employ FaceForensics++ (FF++) [45], which consists of four manipulation methods for the fake data (Deepfakes (DF) [10], FaceSwap (FS) [30], Face2Face (F2F) [50], and NeuralTextures (NT) [49]). It can be noted that, for training, we use only the real videos and generate pseudo-fake data using our synthesized method, SBVs. By default, the c23 version of FF++ is adopted, following recent literature [65, 55, 62, 56, 9]. For further validation, we also evaluate on the following datasets: Celeb-DFv2 (CDF) [35], DeepfakeDetection (DFD) [16], Deepfake Detection Challenge Preview (DFDCP) [13], Deepfake Detection Challenge (DFDC) [14], and WildDeepfake (DFW) [68]. Further details on these datasets are provided in the supplementary material.
@@
Real Fake
Xception [45] 77.9
Face X-ray [32] 94.9
LipForensics [24] 97.1
FTCN [65] 98.3
AltFreezing [55] 98.6
StyleLatentFlows [9] 98.4
Ours (HQ) $\times$ 98.4
Ours (Raw) $\times$ 99.9
Table 4: Generalization to unseen manipulations. AUC (%) comparisons on FF++ [45]. The results for comparison are extracted from the original papers.
@@ 4.2 Comparison with State-of-the-art Methods
CDF2 DFD DFDCP DFDC DFW Avg.
$\times$ $\times$ $\times$ $\times$ 61.5 62.8 59.4 58.5 65.2 61.5
$\times$ $\times$ $\times$ 90.7 95.7 87.9 72.2 70.9 83.5($\uparrow$22.0)
$\times$ $\times$ 91.1 96.0 87.6 72.6 71.0 83.7($\uparrow$22.2)
$\times$ 92.2 95.4 88.5 72.8 71.3 84.0($\uparrow$22.5)
$\times$ $\times$ 93.4 98.5 88.5 72.8 69.6 84.6($\uparrow$23.2)
92.4 98.5 90.0 74.6 74.2 85.9($\uparrow$24.4)
@@
Generalization on Heavily Compressed Data. Following previous work [37, 6], we also evaluate FakeSTormer on heavily compressed (c40) FF++ data. In addition to comparing with several SOTA methods, we train ResNet3D [26], commonly used in deepfake video detection [65, 55, 9], and TimeSformer [2] on NT, then test on DF and FS. The comparison results are presented in Table 3. Our method achieves notably higher AUC scores than other methods across both testing subsets, highlighting its robust generalization capability under various data compression conditions.
@@
CSP Lmk Interpolation Test set AUC (%)
CDF2 DFD DFDCP DFDC DFW Avg.
$\times$ $\times$ 48.4 49.1 48.9 51.7 52.8 50.2
$\times$ 84.1 89.2 86.1 69.5 65.5 78.9($\uparrow$28.7)
90.7 95.7 87.9 72.2 70.9 83.5($\uparrow$33.3)
Table 6: Ablation study of SBVs’ components. Performance analyses of different SBVs’ components using cross-evaluation on multiple datasets [35, 16, 13, 14, 68].
Robustness to Unseen Perturbations. Deepfakes are widely shared on social media, where various perturbations can affect their appearance. Following [28], we evaluate FakeSTormer’s robustness across six unseen degradation types at five levels, comparing it with other augmented-based methods [34, 32, 47, 42]. Figure 5 shows AUC scores for each method on these perturbations, using models trained on FF++. Our results demonstrate that FakeSTormer outperforms prior methods on most distortions, with a slight drop compared to LAA-Net [42] for Change Saturation. Nonetheless, FakeSTormer achieves higher performance on average, especially at higher severity levels, highlighting its superior generalization and robustness. Detailed scores are in the supplementary materials.
@@ 4.3 Additional Discussions
Figure 6: Interpretability. Visualization of spatial and temporal branch outputs on different unseen datasets. The number on each frame denotes the output of the spatial branch $g$, while the heatmap visualizes the output of the temporal branch $h$.
Ablation Study of the FakeSTormer’s Components. We conduct ablation studies to assess the impact of each component in our framework, as shown in Table 5. Using TimeSformer trained on FF++ as the baseline, we experiment with different combinations of components: Self-Blended Videos (SBVs), Vulnerability-driven CutOut (V-CutOut), the spatial branch ($g$), and the temporal branch ($h$). Each component improves performance, with SBVs providing the most significant boost by generating high-quality pseudo-fake data that aids generalization to unseen datasets.
Ablation Study of the SBVs’ Components. SBVs enhance SBI [47] with CSP and LmK Interpolation for robust pseudo-fake generation in video data. Table 6 shows that without these components, simply stacking frame-wise SBI fails to produce consistent temporal features, leading to overfitting on spatial artifacts and poor generalization.
Influence of Number of Frames. Increasing the number of frames $T$ provides more fine-grained temporal information for FakeSTormer. In Table 1, we evaluate $T$ values of $4$, $8$, and $16$. Our results show consistent performance improvement with more frames, confirming our hypothesis. However, increasing $T$ also incurs significant computational cost.
@@
CDF2 DFD DFDCP DFDC DFW Avg.
0.9 1 0.1 89.5 91.0 93.2 71.5 74.7 84.0
0.9 10 0.1 92.5 95.7 86.8 71.7 72.7 83.9
0.9 100 0.1 91.6 98.0 87.3 73.6 70.6 84.2
0.8 100 0.2 92.4 98.5 90.0 74.6 74.2 85.9
0.5 100 0.5 92.4 98.0 88.1 74.5 72.1 85.0
Table 7: Impacts of loss balancing factors. AUC (%) comparisons of FakeSTormer trained with different values of $\lambda_{c}$, $\lambda_{h}$, and $\lambda_{g}$ on cross-dataset setup, demonstrating robustness to varying hyperparameter settings.
Impacts of Loss Balancing Factors. We introduce three hyperparameters, $\lambda_{c}$, $\lambda_{g}$, and $\lambda_{h}$ in Eq (13) to balance the training among the three branches of our framework. In Table 7, we analyze the impact of these hyperparameters using various values. Our results show that the method is robust to a range of hyperparameter values, with the best performance achieved when $\lambda_{c}$, $\lambda_{g}$, and $\lambda_{h}$ are set to $0.8$, $0.2$, and $100$, respectively.
Interpretability. Our spatial and temporal branches output the intensity of the spatial artifacts in each video frame and the vulnerability change over time, respectively. These outputs provide valuable insights during testing, offering interpretability for our model. The spatial branch provides frame-level scores, while the temporal branch offers more fine-grained insights. As shown in Figure 6, the spatial branch output (denoted by the numbers on each frame) gives high values for fake data and low values for real data. For the temporal branch output, the heatmap shows the vulnerability change between each frame and the previous one. As observed, it primarily focuses around the blending boundaries.

### v3

@@ 3.2 FakeSTormer
Our multi-task framework, called FakeSTormer, is inspired by [42, 41], where auxiliary branches are designed to push the feature extractor to focus on vulnerabilities. As discussed earlier, the vulnerability is defined in [42, 41] as the pixels/patches that are the most likely to be impacted by blending artifacts. This strategy is therefore claimed to allow the detection of subtle artifacts that are generic across different types of manipulations. While such a vulnerability-driven approach has shown very promising results [42, 41], it does not take into account the temporal nature of videos. Therefore, in addition to spatial vulnerabilities, we argue that there is a need to model temporal vulnerabilities, which we define as significant temporal changes in the blending boundary. Specifically, we introduce two additional branches, namely a temporal head $h$ and a spatial one $g$. The branch $h$ predicts the derivatives of the blending boundary over time which can reflect high changes, typically characterizing temporal artifacts. Moreover, we suggest the use of a spatial branch $g$ which enables predicting soft labels representing the forgery intensity encoded in each frame, computed from vulnerability information. The proposed framework relies on the TimeSformer backbone [2], which we revisit for better modeling spatial and temporal information.
@@ 3.2.1 Backbone: Revisited TimeSformer
For that purpose, we attach in each dimension of $\mathbf{Z}^{0}$, a spatial token $\mathbf{z}_{s}^{0}\in\mathbb{R}^{D}$ and a temporal token $\mathbf{z}_{t}^{0}\in\mathbb{R}^{D}$, respectively. These tokens will independently interact only with patch embeddings belonging to their dimension axis by leveraging the decomposed SA [2]. This mechanism not only facilitates the disentanglement learning process of spatio-temporal features but is also beneficial to optimize the computational complexity of $\mathcal{O}(T^{2}+N^{2})$ as compared to $\mathcal{O}(T^{2}\cdot N^{2})$ in vanilla SA. Those tokens will be then fed into $L$ ($L=12$ as default) transformer encoder blocks, as described in Figure 2-I. Formally, the feature extraction process can be summarized as follows,
@@ 3.2.2 Temporal Head $h$
Ground Truths. Our temporal head $h$ aims to model fine-grained temporal vulnerabilities in deepfake videos through a regression task. First, to generate ground truth data for the branch $h$, we hypothesize that temporal high-changes in the blending boundary can reflect the presence of temporal artifacts (see Figure 2-IV). To achieve this, we compute $\mathbf{D}$ based on $\tilde{\mathbf{B}}$ such that:
@@
Objective Function. For training the temporal branch, we optimize the following Mean Squared Error (MSE) loss,
@@ 3.2.3 Spatial Head $g$
Ground Truths. To avoid overfitting one type of artifact, we enforce the model to explicitly predict soft labels representing the intensity level of spatial artifacts for each video frame. Note that several works [4, 52, 49, 66, 40] have leveraged soft labels for training regularization. Given a pseudo-fake video $\mathbf{X}=(\mathbf{X}(t))_{t\in[[1,T]]}$ formed by $T$ frames and $\tilde{\mathbf{B}}=(\tilde{\mathbf{B}}(t))_{t\in[[1,T]]}$ its associated cutout blending boundary, the ground truth for these soft labels is generated for each frame $t$ as follows,
@@ 3.2.4 Overall Training Objective
Method Training Test set AUC (%)
Real Fake CDF DFD DFDCP DFDC DFW DiffSwap
Xception [46] 73.7 - - 70.9 - -
MATT [69] 68.3 92.9 63.0 - 65.7 -
RECCE [3] 70.9 98.2 - - 68.2 -
SBI [48] $\times$ 90.6 - - 72.4 - -
SFDG [55] 75.8 88.0 73.6 - 69.3 -
LSDA [60] 91.1 - 77.0 - - -
STIL [23] 75.6 - - - - -
LipForensics [25] 82.4 - - 73.5 - -
RealForensics [26] 86.9 82.2 75.9 - - -
FTCN [71] 86.9 94.4 74.0 71.0 - -
ISTVT [67] 84.1 - 74.2 - - -
AltFreezing [56] 89.5 98.5 - - - -
Swin+TALL [57] 90.8 - 76.8 - - -
StyleLatentFlows [9] 89.0 96.1 - - - -
LFGDIN [63] 90.4 - 80.8 - - 85.7
FakeSTormer ($T=4$) $\times$ 92.4 98.5 90.0 74.6 74.2 96.9
FakeSTormer ($T=8$) $\times$ 92.4 98.2 90.0 74.9 75.9 97.1
FakeSTormer ($T=16$) $\times$ 92.8 98.6 90.2 75.1 75.3 97.2
Table 1: Generalization to unseen datasets. AUC (%) comparisons at video-level on multiple unseen datasets [36, 17, 13, 14, 74, 70]. All detectors are trained on FF++(c23). Results are directly extracted from the original papers and from [42, 25]. Bold and Underlined text, respectively highlight the best and the second best performance, excluding the variants of our framework with $T=8$ and $T=16$.
@@ 4 Experiment
Method Training set Cross-dataset DF40 subset
Real Fake CDF DFDCP DiffSwap BlendFace FSGAN MobileSwap
Face X-ray [33] 79.5 - - - - -
PCL+I2G [68] 90.0 74.3 - - - -
SLADD [6] 79.7 - - - - -
SBI [48] $\times$ 93.2 86.2 90.6 86.5 85.4 86.6
LAA-Net [42] $\times$ 95.4 86.9 92.1 91.2 94.2 93.9
STC-Scratch [34] $\times$ 83.4 86.8 - - - -
STC-Pretrain [34] $\times$ 95.8 89.4 - - - -
ST-SBV [24] $\times$ 90.3 91.2 - - - -
StA+VB [62] $\times$ 94.7 90.9 - 90.6 96.4 94.6
TimeSformer [2] + SBV $\times$ 94.9 93.0 93.3 89.7 94.6 94.6
FakeSTormer $\times$ 96.5 94.1 97.7 91.1 96.4 95.0
Table 2: AUC(%) comparison at video-level with other data synthesis methods. For fair comparison, we train our FakeSTormer on raw data of FF++(c0), and test cross-dataset on [36, 13, 70] and cross-manipulation on three subsets of [61].
@@ 4.1 Settings
Datasets. We set up our datasets following several works [56, 71, 57, 67, 9, 59, 55, 3]. For both training and validation, we employ FaceForensics++ (FF++) [46], which consists of four manipulation methods for the fake data (Deepfakes (DF) [10], FaceSwap (FS) [31], Face2Face (F2F) [51], and NeuralTextures (NT) [50]). It can be noted that, for training, we use only the real videos and generate pseudo-fake data using our synthesized method, SBV. By default, the c23 version of FF++ is adopted, following the recent literature [71, 56, 67, 57, 9]. For further validation, we also evaluate on the following datasets: Celeb-DFv2 (CDF) [36], DeepfakeDetection (DFD) [17], Deepfake Detection Challenge Preview (DFDCP) [13], Deepfake Detection Challenge (DFDC) [14], WildDeepfake (DFW) [74], DF40 [61], and DiffSwap [63, 70] generated using a recent diffusion-based approach [70]. Further details on these datasets are provided in supplementary materials.
@@
Real Fake DF FS F2F NT Avg.
Xception [46] 93.9 51.2 86.8 79.7 77.9
Face X-ray [33] 99.5 93.2 94.5 92.5 94.9
SBI [48] $\times$ 98.6 95.4 92.6 82.3 92.2
LSDA [60] 96.9 95.1 96.4 94.9 95.8
LipForensics [25] 99.7 90.1 99.7 99.1 97.1
FTCN [71] 99.8 99.6 98.2 95.6 98.3
RealForensics [26] 100 97.1 99.7 99.2 99.0
AltFreezing [56] 99.8 99.7 98.6 96.2 98.6
StyleLatentFlows [9] 99.7 98.8 98.6 96.4 98.4
NACO [65] 99.9 99.7 99.8 99.4 99.7
LFGDIN [63] 96.2 80.5 90.5 81.7 87.2
Ours (c23) $\times$ 99.9 97.8 98.5 97.2 98.4
Ours (c0) $\times$ 100 99.8 99.9 99.7 99.9
Table 4: Generalization to unseen manipulations. AUC (%) comparisons on FF++ [46], which consists of four manipulation methods (DF, FS, F2F, NT).
@@ 4.2 Comparison with State-of-the-art Methods
CDF DFD DFDCP DFDC DFW DiffSwap Avg.
$\times$ $\times$ $\times$ $\times$ 61.5 62.8 59.4 58.5 65.2 71.6 63.2
$\times$ $\times$ $\times$ 90.7 95.7 87.9 72.2 70.9 92.9 85.1($\uparrow$21.9)
$\times$ $\times$ 91.1 96.0 87.6 72.6 71.0 93.1 85.2($\uparrow$22.0)
$\times$ 92.2 95.4 88.5 72.8 71.3 93.8 85.7($\uparrow$22.5)
$\times$ $\times$ 93.4 98.5 88.5 72.8 69.6 97.3 86.7($\uparrow$23.5)
92.4 98.5 90.0 74.6 74.2 96.9 87.8($\uparrow$24.6)
@@
Generalization on Heavily Compressed Data. Following previous work [38, 6], we also evaluate FakeSTormer on heavily compressed FF++(c40) data. In addition to comparing with several SOTA methods, we train ResNet3D [27], commonly used in deepfake video detection [71, 56, 9], and TimeSformer [2] on NT, then test on DF and FS. The comparison results are presented in Table 3. Our method achieves notably higher AUC scores than other methods across both testing subsets, highlighting its robust generalization capability under various data compression conditions.
@@
Robustness to Unseen Perturbations. Deepfakes are widely shared on social media, where various perturbations can affect their appearance. Following [29], we evaluate FakeSTormer’s robustness across six unseen degradation types at five levels, comparing it with other augmented-based methods [35, 33, 48, 42]. Figure 3 shows AUC scores for each method on these perturbations, using models trained on FF++. Our results demonstrate that FakeSTormer outperforms prior methods on most distortions, with a slight drop compared to LAA-Net [42] for Change Saturation. Nonetheless, FakeSTormer achieves higher performance on average, especially at higher severity levels, highlighting its superior generalization and robustness. Detailed scores are in supplementary materials.
@@ 4.3 Additional Discussions
Ablation Study of the FakeSTormer’s Components. We conduct ablation studies to assess the impact of each component in our framework, as shown in Table 5. Using TimeSformer trained on FF++ as the baseline, we experiment with different combinations of components: Self-Blended Video (SBV), Vulnerability-driven CutOut (V-CutOut), the spatial branch ($g$), and the temporal branch ($h$). Each component improves performance, with SBV providing the most significant boost by generating high-quality pseudo-fake data that aids generalization to unseen datasets.
Ablation Study of the SBV’s Components. SBV enhances SBI [48] with CSP and LI for robust pseudo-fake generation in video data. Table 6 shows that without these components, simply stacking frame-wise SBIs fails to produce consistent temporal features, leading to overfitting on more obvious artifacts and poor generalization [56]. A qualitative comparison is provided in supplementary materials.
Method Comp. CSP LI Test set AUC (%)
CDF DFD DFDCP DFDC DFW DiffSwap
Stacked SBIs c23 $\times$ $\times$ 48.4 49.1 48.9 51.7 52.8 55.5
+ CSP c23 $\times$ 84.1 89.2 86.1 69.5 65.5 85.7
SBV c23 90.7 95.7 87.9 72.2 70.9 92.9
SBV c0 94.9 97.6 93.0 76.4 75.3 93.3
Table 6: Ablation study of SBV’s components. Performance analyses of different SBV’s components using cross-evaluation on multiple datasets [36, 17, 13, 14, 74, 70].
Influence of Number of Frames. Increasing the number of frames $T$ provides more fine-grained temporal information. In Table 1, we vary $T$ values by fixing it to $4$, $8$, and $16$. Our results show a consistent performance improvement with more frames, confirming our hypothesis. However, increasing $T$ also incurs a higher computational cost.
@@
CDF DFD DFDCP DFDC DFW DiffSwap Avg.
0.9 1 0.1 89.5 91.0 93.2 71.5 74.7 90.6 85.1
0.9 10 0.1 92.5 95.7 86.8 71.7 72.7 94.5 85.7
0.9 100 0.1 91.6 98.0 87.3 73.6 70.6 96.2 86.2
0.8 100 0.2 92.4 98.5 90.0 74.6 74.2 96.9 87.8
0.5 100 0.5 92.4 98.0 88.1 74.5 72.1 97.1 87.0
Table 7: Impact of loss balancing factors. AUC (%) comparisons of FakeSTormer trained with different values of $\lambda_{c}$, $\lambda_{h}$, and $\lambda_{g}$ on cross-dataset setup, demonstrating robustness to varying hyperparameter settings.
Impact of Loss Balancing Factors. We introduce three hyperparameters, $\lambda_{c}$, $\lambda_{g}$, and $\lambda_{h}$ in Eq (13) to balance the training among the three branches of our framework. In Table 7, we analyze the impact of these hyperparameters using various values. Our results show that the method is robust to a range of hyperparameter values, with the best performance achieved when $\lambda_{c}$, $\lambda_{g}$, and $\lambda_{h}$ are set to $0.8$, $0.2$, and $100$, respectively.
4.4 Visualization of Saliency Maps